# Simulating a two component Bose-Hubbard model with imbalanced hopping in a Rydberg tweezer array

### v1

@@ Abstract
Optical tweezer arrays of neutral atoms provide a versatile platform for quantum simulation due to the range of interactions and Hamiltonians that can be realized and explored. We propose to simulate a two-component Bose-Hubbard model with power-law hopping using arrays of multilevel Rydberg atoms featuring resonant dipolar interactions. The diversity of states that can be used to encode the local Hilbert space of the Bose-Hubbard model enables control of the relative hopping rate of each component and even the realization of spin-flip hopping. We use numerical simulations to show how multilevel Rydberg atoms provide an opportunity to explore the diverse non-equilibrium quench dynamics of the model. For example, we demonstrate a separation of the relaxation timescales of effective spin and charge degrees of freedom, and observe regimes of slow relaxation when the effective hopping rates of the two components are vastly different due to dynamical constraints arising from hardcore boson interactions. We discuss the technical details of realizing our proposal in state-of-the-art Rydberg tweezer arrays.
@@ I Introduction
Programmable tweezer arrays of Rydberg atoms provide a versatile platform for quantum simulation of many models of quantum magnetism in a range of geometries Morgado and Whitlock (2021); Kaufman and Ni (2021); Barredo et al. (2018); Weimer et al. (2010); Morgado and Whitlock (2021); Labuhn et al. (2016a); De Léséleuc et al. (2019) and present exciting prospects for realizing exotic quantum phases of matter Verresen et al. (2021); Semeghini et al. (2021); Samajdar et al. (2021). By encoding a qubit in Rydberg and/or ground states of an atom, one can selectively realize $1/r^{6}$ van der Waals or anisotropic $1/r^{3}$ dipolar interactions that naturally emulate Ising Schauss (2018); Labuhn et al. (2016b); Borish et al. (2020); Bernien et al. (2017); Song et al. (2021); Eckner et al. (2023), XY Bornet et al. (2023); Orioli et al. (2018) or XXZ Signoles et al. (2021) spin models. Recent work has demonstrated further versatility with applied fields that realize local or global driving and energy shifts, which can be exploited to realize Floquet Hamiltonians Geier et al. (2021); Scholl et al. (2022); Christakis et al. (2023); Nishad et al. (2023). These tools for Hamiltonian engineering are combined with technical capabilities such as site-resolved imaging and state-preparation in the tweezer apparatus.
@@
New opportunities are emerging in these systems by applying the aforementioned levels of control to multiple (i.e., more than two) Rydberg levels within each atom. In general, multilevel atoms present a promising platform to expand the frontiers of quantum simulation beyond, e.g., well-established spin-$1/2$ models, and controllably enrich the complexity of many-body dynamics. Recent work involving Rydberg atoms has demonstrated how internal structure can be exploited to, e.g., realize synthetic dimensions Ozawa and Price (2019); Hazzard and Gadway (2023). For example, a set of neighbouring Rydberg states can be coupled by independent microwaves to study topological physics Kanungo et al. (2022) or non-equilibrium dynamics featuring strong interactions Chen et al. (2023a). In this work, we propose to use resonant dipolar exchange interactions between three Rydberg states with alternating parity to simulate the non-equilibrium dynamics of a two component (i.e., spin-$1/2$) Bose-Hubbard model featuring tunable power-law hopping and hardcore interactions. Single atoms are trapped in a programmable tweezer array, realizing the spatial dimension, while the Rydberg levels encode the local Hilbert space of the model, corresponding to the onsite occupation of each boson component. Our mapping explicitly excludes double occupancy of any site to realize perfect hardcore interactions between the effective bosons, while the natural variation of exchange interactions between different levels allows us to easily investigate different hopping rates for each component or even resonant spin-flip tunneling.
@@ II Quantum simulation with multilevel Rydberg atoms
$\hat{H}$ $=\sum_{m=1}^{n-1}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\mathcal{N}}J^{m}_{ij}\left(|m+1\rangle\langle m|\right)_{i}\left(|m\rangle\langle m+1|\right)_{j}$ (3)
@@
The first encoding of the boson occupancies is via the mapping |$\ket{1}_{j}\leftrightarrow\ket{\uparrow}_{j}$, |$\ket{2}_{j}\leftrightarrow\ket{0}_{j}$ and |$\ket{3}_{j}\leftrightarrow\ket{\downarrow}_{j}$, where |$\ket{\uparrow}_{j}$ (|$\ket{\downarrow}_{j}$) corresponds to the site $j$ being occupied with an $\uparrow$ ($\downarrow$) boson and |$\ket{0}_{j}$ encoding an empty site [see Fig. 1(b)]. Then, Eq. (3) can be written as a two-component Bose-Hubbard Hamiltonian with component-dependent hopping,
@@
Here, $\hat{b}_{\sigma j}^{\dagger}$($\hat{b}_{\sigma j}$) is the creation (annihilation) operator of a spin $\sigma$ boson at site $j$, with $\sigma=\uparrow$ or $\downarrow$. The hardcore interaction enters through the requirement that $\hat{b}_{\sigma j}^{\dagger}\hat{b}_{\sigma^{\prime}j}^{\dagger}=\hat{b}_{\sigma j}\hat{b}_{\sigma^{\prime}j}=0$. The first line of Eq. (5) describes hopping of each component between sites $i$ and $j$ with rates $J^{1}_{ij}$ and $J^{2}_{ij}$, respectively, which can be controlled by the choice of Rydberg levels used in the simulator. In the absence of the microwave terms in the second line, which lead to the coherent creation/destruction of bosons at each site, the Hamiltonian separately conserves the total number of $\uparrow$ and $\downarrow$ bosons on the lattice, $N_{\uparrow}=\sum_{i}\hat{b}^{\dagger}_{\uparrow i}\hat{b}_{\uparrow i}$ and $N_{\downarrow}=\sum_{i}\hat{b}^{\dagger}_{\downarrow i}\hat{b}_{\downarrow i}$, respectively. This can be understood as a direct result of the conservation of total population in each Rydberg state under the dipolar exchange interaction.
@@ V Experimental considerations
The physics discussed in the prior sections are accessible with the current capabilities of quantum simulators based on arrays of Rydberg atoms. For example, the preparation of arbitrary product states with site-resolved structure (such as Neel spin ordering or density waves) can be achieved by exploiting light shifts of the initially trapped ground state or Rydberg levels Chen et al. (2023b), combined with independent coherent driving of each Rydberg-Rydberg transition with tunable microwave fields Chen et al. (2023a); Sundar et al. (2019). This coherent control of each atom, combined with site-resolved imaging, also enables the extraction of many-body correlations and the construction of, e.g., number entropies. We note that the extraction of charge observables only requires the measurement of population in a single Rydberg level, which can be achieved by standard detection techniques that transfer population of the targeted Rydberg level back to the ground-state for recapture in the tweezers and subsequent imaging Jau et al. (2016). On the other hand, for spin observables the measurement of populations in all three Rydberg levels could be achieved by selective transfer and shelving of population of two of the Rydberg states into different ground-state magnetic sublevels Lee et al. (2017), with the population of the third state inferred through atom number conservation.
Observing the relaxation dynamics over long timescales will require the identification of appropriate Rydberg levels with sufficiently strong dipolar exchange interactions compared to relevant technical and fundamental limitations on the total simulation time. For example, based on the micro-Kelvin temperatures routinely achieved in tweezer experiments, the absolute simulation time is typically limited to $t\lesssim 20~{}\mu$s to ensure the recapture probability is sufficient for high-fidelity detection (the tweezers are turned off during the simulation time to avoid perturbation of the Rydberg states). This is typically much shorter than the characteristic lifetime of the Rydberg states at room temperature. Second, besides the dipolar terms there can be additional short-range van der Waals interactions $J_{\mathrm{vdW}}/r_{ij}^{6}$ that we have neglected in our model Morgado and Whitlock (2021). These can be minimized by choosing an appropriate tweezer spacing (as the $1/r_{ij}^{6}$ dependence will fall off much faster than the $1/r_{ij}^{3}$ dipolar terms), although the details are highly dependent on the specific choice of atomic species and Rydberg states. Additional off-resonant dipolar interactions to other Rydberg states can be suppressed by choosing an appropriately strong quantization field Lee et al. (2017).
@@ VI Summary and outlook
Our work demonstrates the opportunities for quantum simulation and studies of non-equilibrium dynamics with multilevel Rydberg atoms. In our example using three internal states, we showed the tunability of the platform can enable controllable studies of a two-component Bose-Hubbard model with tunable, power-law hopping and hardcore interactions, which features rich relaxation dynamics that depend on the interplay of these ingredients. Beyond prior work studying dynamical constraints, localization and thermalization in this and similar models Gadway et al. (2011); Sarkar et al. (2020); Lan et al. (2018); Schiulaz et al. (2015); Yao et al. (2016); Sirker (2019); Darkwah Oppong et al. (2022), we also highlight potential connections to recent studies of glassy dynamics in ensembles of Rydberg qubits featuring positional disorder Signoles et al. (2021). For instance, preparing an initial product state where one component is immobile (i.e., $J_{1}/J_{2}=0$) and is randomly scattered to populate a variable fraction of the chain while the remaining sites are in a superposition of empty and occupied by the mobile component, our model directly maps to typically studied configurations. An interesting direction would be to perturb away from this scenario by allowing the previously immobile component to slowly move ($J_{1}/J_{2}\ll 1$). Note that the predictions we discuss in this work can also be realized in other systems featuring multilevel internal structure and exchange interactions, such as polar molecules Sundar et al. (2018, 2019).
The versatility and breadth of the Rydberg platform also suggests a number of interesting extensions for future work. For example, while we have already demonstrated that the choice of basis states in the three-level system can be used to generate distinct spin-flip hopping, working at small tweezer spacings to purposely strengthen the van der Waals interactions between Rydberg atoms can be used to engineer additional $1/r^{6}$ density-density interactions in the Bose-Hubbard model that can compete with the $1/r^{3}$ hopping. Microwave driving of the different Rydberg transitions can also enable studies of Floquet-engineered Hamiltonians in the spirit of recent work with Rydberg qubits and polar molecules Geier et al. (2021); Scholl et al. (2022); Christakis et al. (2023); Nishad et al. (2023). Lastly, all of these tools can similarly be applied to study non-equilibrium dynamics in higher-dimensional arrays and ladders of more than three Rydberg states, although for the latter a description in terms of hardcore bosons is no longer readily applicable.

### v2

@@ Abstract
Optical tweezer arrays of neutral atoms provide a versatile platform for quantum simulation due to the range of interactions and Hamiltonians that can be realized and explored. We propose to simulate a two-component Bose-Hubbard model with power-law hopping using arrays of multilevel Rydberg atoms featuring resonant dipolar interactions. The diversity of states that can be used to encode the local Hilbert space of the Bose-Hubbard model enables control of the relative hopping rate of each component and even the realization of spin-flip hopping. We use numerical simulations to show how multilevel Rydberg atoms provide an opportunity to explore the diverse non-equilibrium quench dynamics of the model. For example, we demonstrate a separation of the relaxation timescales of effective spin and charge degrees of freedom, and observe regimes of slow relaxation when the effective hopping rates of the two components are vastly different due to dynamical constraints arising from hardcore boson interactions. We discuss prospects for studying these effects in state-of-the-art Rydberg tweezer arrays.
@@ I Introduction
Programmable tweezer arrays of Rydberg atoms provide a versatile platform for quantum simulation of many models of quantum magnetism in a range of geometries Morgado and Whitlock (2021); Kaufman and Ni (2021); Barredo et al. (2018); Weimer et al. (2010); Morgado and Whitlock (2021); Labuhn et al. (2016a); De Léséleuc et al. (2019) and present exciting prospects for realizing exotic quantum phases of matter Verresen et al. (2021); Semeghini et al. (2021); Samajdar et al. (2021). By encoding a qubit in Rydberg and/or ground states of an atom, one can selectively realize $1/r^{6}$ van der Waals or anisotropic $1/r^{3}$ dipolar interactions that naturally emulate Ising Schauss (2018); Labuhn et al. (2016b); Borish et al. (2020); Bernien et al. (2017); Song et al. (2021); Eckner et al. (2023), XY Bornet et al. (2023); Orioli et al. (2018), or XXZ Signoles et al. (2021) spin models. Recent work has demonstrated further versatility with applied fields that realize local or global driving and energy shifts, which can be exploited to realize Floquet Hamiltonians Geier et al. (2021); Scholl et al. (2022); Christakis et al. (2023); Nishad et al. (2023). These tools for Hamiltonian engineering are combined with technical capabilities such as site-resolved imaging and state-preparation in the tweezer apparatus.
@@
New opportunities are emerging in these systems by applying the aforementioned levels of control to multiple (i.e., more than two) Rydberg levels within each atom. In general, multilevel atoms present a promising platform to expand the frontiers of quantum simulation beyond, e.g., well-established spin-$1/2$ models, and controllably enrich the complexity of many-body dynamics. Recent work involving Rydberg atoms has demonstrated how internal structure can be exploited to, e.g., realize synthetic dimensions Ozawa and Price (2019); Hazzard and Gadway (2023). For example, a set of neighbouring Rydberg states can be coupled by independent microwaves to study topological physics Kanungo et al. (2022), non-equilibrium dynamics featuring strong interactions Chen et al. (2023a), or classical analog of quantum models Cohen et al. (2024). In this work, we propose to use resonant dipolar exchange interactions between three Rydberg states with alternating parity to simulate the non-equilibrium dynamics of a two component (i.e., spin-$1/2$) Bose-Hubbard model featuring tunable power-law hopping and hardcore interactions. Single atoms are trapped in a programmable tweezer array, realizing the spatial dimension, while the Rydberg levels encode the local Hilbert space of the model, corresponding to the onsite occupation of each boson component. Our mapping explicitly excludes double occupancy of any site to realize perfect hardcore interactions between the effective bosons, while the natural variation of exchange interactions between different levels allows us to easily investigate different hopping rates for each component or even resonant spin-flip tunneling. We note that multilevel Rydberg atoms have also similarly been proposed for quantum simulation of a related Bosonic $t-J$ model in a recent work Homeier et al. (2023).
@@ II Quantum simulation with multilevel Rydberg atoms
$\hat{H}$ $=\sum_{m=1}^{n-1}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\mathcal{N}}J^{m}_{ij}\left(|m+1\rangle\langle m|\right)_{i}\left(|m\rangle\langle m+1|\right)_{j}$
@@
The first encoding of the boson occupancies is via the mapping |$\ket{1}_{j}\leftrightarrow\ket{\uparrow}_{j}$, |$\ket{2}_{j}\leftrightarrow\ket{0}_{j}$ and |$\ket{3}_{j}\leftrightarrow\ket{\downarrow}_{j}$, where |$\ket{\uparrow}_{j}$ (|$\ket{\downarrow}_{j}$) corresponds to the site $j$ being occupied with an $\uparrow$ ($\downarrow$) boson and |$\ket{0}_{j}$ encoding an empty site [see Fig. 1(b)]. Then, Eq. (II) can be written as a two-component Bose-Hubbard Hamiltonian with component-dependent hopping,
@@
Here, $\hat{b}_{\sigma j}^{\dagger}$($\hat{b}_{\sigma j}$) is the creation (annihilation) operator of a spin $\sigma$ boson at site $j$, with $\sigma=\uparrow$ or $\downarrow$. The hardcore interaction enters through the requirement that $\hat{b}_{\sigma j}^{\dagger}\hat{b}_{\sigma^{\prime}j}^{\dagger}=\hat{b}_{\sigma j}\hat{b}_{\sigma^{\prime}j}=0$. The first line of Eq. (2) describes hopping of each component between sites $i$ and $j$ with rates $J^{1}_{ij}$ and $J^{2}_{ij}$, respectively, which can be controlled by the choice of Rydberg levels used in the simulator. In the absence of the microwave terms in the second line, which lead to the coherent creation/destruction of bosons at each site, the Hamiltonian separately conserves the total number of $\uparrow$ and $\downarrow$ bosons on the lattice, $N_{\uparrow}=\sum_{i}\hat{b}^{\dagger}_{\uparrow i}\hat{b}_{\uparrow i}$ and $N_{\downarrow}=\sum_{i}\hat{b}^{\dagger}_{\downarrow i}\hat{b}_{\downarrow i}$, respectively. This can be understood as a direct result of the conservation of total population in each Rydberg state under the dipolar exchange interaction. Note that in the recent work of Ref. Homeier et al. (2023), it was proposed to similarly use three Rydberg levels to emulate a bosonic $t-J$ model with hardcore constraints. Our Hamiltonian (2) with $J_{ij}^{1}=J_{ij}^{2}$ and $\Omega_{1}=\Omega_{2}=0$ corresponds to a limiting case of that work.
@@ V Experimental considerations
The physics discussed in the prior sections is within reach of state-of-the-art quantum simulators based on tweezer arrays of Rydberg atoms. For example, the preparation of arbitrary product states with site-resolved structure (such as Neel spin ordering or density waves) can be achieved by exploiting light shifts of the initially trapped ground state or Rydberg levels Chen et al. (2023b), combined with independent coherent driving of each Rydberg-Rydberg transition with tunable microwave fields Chen et al. (2023a); Sundar et al. (2019). This coherent control of each atom, combined with high fidelity site-resolved imaging Ebadi et al. (2021); Madjarov et al. (2020); Evered et al. (2023); Bornet et al. (2024); Manetsch et al. (2024), can enable the extraction of many-body correlations and the construction of, e.g., number entropies. We note that the extraction of charge observables only requires the measurement of population in a single Rydberg level, which can be achieved by standard detection techniques that transfer population of the targeted Rydberg level back to the ground-state for recapture in the tweezers and subsequent imaging Jau et al. (2016). For spin observables, one can in principle measure the populations in all three Rydberg levels by selective transfer and shelving of population of each Rydberg state into different ground-state magnetic sublevels Lee et al. (2017).
Observing the relaxation dynamics over long timescales requires the identification of Rydberg levels with sufficiently strong dipolar exchange interactions compared to relevant technical and fundamental limitations on the total simulation time. One such constraint is residual thermal motion of the atoms that limits recapture of the atoms by the tweezers for subsequent imaging (the tweezers are turned off during the quench dynamics to avoid perturbation of the Rydberg states). For the micro-Kelvin temperatures routinely achieved in state-of-the-art tweezer experiments this leads to a first limit of about $t\lesssim 20~{}\mu$s. A second constraint is the finite lifetime of the chosen Rydberg states. While the effects of this can in principle be included in numerical simulations comparing to a specific experimental apparatus, this is beyond the scope of the current work. Instead, we highlight that any effects can be suppressed by a combination of state-resolved single-shot detection (see above) and post-selection. In particular, state-resolved detection of all three Rydberg states in a single experimental shot enables discrimination of cases where an atom decays out of the targeted Rydberg levels during the quench dynamics. This approach is feasible as Rydberg quantum simulators typically have relatively fast duty cycles and the fraction of experimental shots where no loss occurs can be sufficiently large for the moderate sized arrays discussed in this work. The latter is approximately quantified by the survival probability for $\mathcal{N}$ atoms, $P_{\mathcal{N}}=e^{-\mathcal{N}t/\tau}$, where $t$ is the quench duration and $\tau$ is a characteristic lifetime of the relevant Rydberg states.
As an illustrative example, Table 1 shows relevant parameters for candidate Rydberg levels in cesium. The states and related parameters are chosen by requiring that timescales of the relevant physics (motivated by Figs. 4 and 6) can be reached with a survival probability of at least $P_{\mathcal{N}}\geq 0.1$ with $\mathcal{N}=10-20$, as well as other factors such as feasible tweezer separation, reasonable simulation time (i.e., sufficient recapture probability as discussed above) and suppressed additional short-range van der Waals interactions $J_{\mathrm{vdW}}/r_{ij}^{6}$ neglected in our model Morgado and Whitlock (2021). Moreover, additional off-resonant dipolar interactions to other Rydberg state magnetic sublevels can be feasibly suppressed with an appropriately strong quantization field Lee et al. (2017); Evered et al. (2023). In general, a far greater diversity of possible level configurations can be explored, including by consideration of other atomic species, with an eye to tuning the precise values of $J_{1,2}$ (and associated parameters) or optimizing for other regimes (i.e., shorter timescales such as those probed in Fig. 2).
Configuration State Lifetime $\left(\rm\mu s\right)$ Level $C_{3}$/h${}\left(\rm MHz\cdot\mu m^{3}\right)$ Spacing $\left(\rm\mu m\right)$ $J_{i}/2\pi~{}(\rm MHz)$
$|1\rangle$ 51 51F5/2,mj=1/2 379, ($|1\rangle\leftrightarrow|2\rangle$) $J_{1}/{2\pi}$ = 0.74
$J_{1}/J_{2}=0.082$ $|2\rangle$ 55 53D3/2,mj=3/2 4625, $\left(|2\rangle\leftrightarrow|3\rangle\right)$ 8 $J_{2}/{2\pi}$ = 9.03
$|3\rangle$ 121 54P1/2,mj=1/2
$|1\rangle$ 51 51F7/2,mj=3/2 1973, ($|1\rangle\leftrightarrow|2\rangle$) $J_{1}/{2\pi}$ = 2.71
$J_{1}/J_{2}=0.877$ $|2\rangle$ 55 53D5/2,mj=3/2 2246, ($|2\rangle\leftrightarrow|3\rangle$) 9 $J_{2}/{2\pi}$ = 3.08
$|3\rangle$ 118 54P3/2,mj=3/2
$|1\rangle$ 84 62D5/2,mj=1/2 3244.8, ($|1\rangle\leftrightarrow|2\rangle$) $J_{1}/{2\pi}$ = 0.7922
$J_{1}/J_{2}=1.004$ $|2\rangle$ 169 63P3/2,mj=1/2 3232.3, ($|2\rangle\leftrightarrow|3\rangle$) 16 $J_{2}/{2\pi}$ = 0.7891
$|3\rangle$ 112 64S1/2,mj=1/2
Table 1: Summary of candidate Rydberg states to achieve target values of $J_{1}/J_{2}$ibalić et al. (2017). For the case of $J_{1}/J_{2}=0.082$ the total survival probability for the atom array is in the range of about $0.70.5<P_{\mathcal{N}}<$ 0.7 for $10\leq\mathcal{N}\leq 20$ at $J_{2}t=100$. Similarly, for $J_{1}/J_{2}=0.877$ we have $0.360.13<P_{\mathcal{N}}<$ 0.36 at $J_{2}t=100$, and for $J_{1}/J_{2}=1.004$ we have $0.490.24<P_{\mathcal{N}}<$ 0.49 at $J_{2}t=30$, both using the same range of array sizes.
@@ VI Summary and outlook
Our work demonstrates the opportunities for quantum simulation and studies of non-equilibrium dynamics with multilevel Rydberg atoms. In our example using three internal states, we showed the tunability of the platform can enable controllable studies of a two-component Bose-Hubbard model with tunable, power-law hopping and hardcore interactions, which features rich relaxation dynamics that depend on the interplay of these ingredients. Beyond prior work studying dynamical constraints, localization and thermalization in this and similar models Gadway et al. (2011); Sarkar et al. (2020); Lan et al. (2018); Schiulaz et al. (2015); Yao et al. (2016); Sirker (2019); Darkwah Oppong et al. (2022), we also highlight potential connections to recent studies of glassy dynamics in ensembles of Rydberg qubits featuring positional disorder Signoles et al. (2021). For instance, preparing an initial product state where one component is immobile (i.e., $J_{1}/J_{2}=0$) and is randomly scattered to populate a variable fraction of the chain while the remaining sites are in a superposition of empty and occupied by the mobile component, our model directly maps to typically studied configurations. An interesting direction would be to perturb away from this scenario by allowing the previously immobile component to slowly move ($J_{1}/J_{2}\ll 1$). Note that the predictions we discuss in this work can also be realized in other systems featuring multilevel internal structure and exchange interactions, such as polar molecules Sundar et al. (2018, 2019); Cornish et al. (2024).
The versatility and breadth of the Rydberg platform also suggests a number of interesting extensions for future work. For example, while we have already demonstrated that the choice of basis states in the three-level system can be used to generate distinct spin-flip hopping, working at small tweezer spacings to purposely strengthen the van der Waals interactions between Rydberg atoms can be used to engineer additional $1/r^{6}$ density-density interactions in the Bose-Hubbard model that can compete with the $1/r^{3}$ hopping. Microwave driving of the different Rydberg transitions can also enable studies of Floquet-engineered Hamiltonians in the spirit of recent work with Rydberg qubits and polar molecules Geier et al. (2021); Scholl et al. (2022); Christakis et al. (2023); Nishad et al. (2023); Homeier et al. (2023). Lastly, all of these tools can similarly be applied to study non-equilibrium dynamics in higher-dimensional arrays and ladders of more than three Rydberg states, although for the latter a description in terms of hardcore bosons is no longer readily applicable.